OpenStudy (anonymous):

Find an equation of the line that satisfies the given: (-3,-5); slope -7/2

OpenStudy (anonymous):

y=-3.5*(x+3)-5 or 3.5x+y+15.5=0

OpenStudy (anonymous):

7x+2y+31=0 looks fine as well.

OpenStudy (anonymous):

how did you get that?

OpenStudy (mimi_x3):

y = mx + b -5 = -7/2(-3) + b -5 = 21/2 + b b = -31/2 y = -7/2x -31/2

OpenStudy (anonymous):

given point (a,b), slope k the equation should be: y=k*(x-a)+b

OpenStudy (anonymous):

ok, i get it..! thanks